A lorry with mass 3500 kg is parked at the top of a steep hill with a 1– in– 8 gradient when its handbrake fails. Assume that the lorry has a constant frictional resistance to motion of 500 N. The lorry rolls 40 m down the hill before crashing into a lamp post. Use the principle of conservation of energy to calculate the velocity of the lorry immediately prior to its impact with the lamp post.


1
Expert's answer
2021-11-25T01:41:40-0500

Acceleration :

"a = \frac{F}{m}= \frac{500}{3500} =0.143 \frac{m}{sec^2}"

velocity of the lorry

"V^2 = u^2+2as"

"V^2= 0+2\times0.143\times 40"

"V=3.38 \frac{m}{sec^2}"